OpenStudy (anonymous):

Verify the identity. cotx minus pi divided by two. = -tan x written as cot(x - (pi/2) = - tan x

OpenStudy (anonymous):

Please check my answer: cot(x - pi/2) = -tan (x) cot(x + 2pi + (3pi/2)) = -tan(x) -tan(x) = -tan(x)

OpenStudy (shinalcantara):

oh dear. what have you done? you're answer is defying the rules of trigonometry..

OpenStudy (shinalcantara):



OpenStudy (shinalcantara):